# On the Pitfalls of Batch Normalization for End-to-End Video Learning: A Study on Surgical Workflow Analysis

### v4

@@ 4.2 Surgical Phase Recognition
CATARACTS: Since CATARACTS contains shorter videos with more fine-grained annotations, the benefits of our context-maximizing strategies are expectedly smaller here. Nevertheless, we can still reproduce several findings (Table 5): Single-sequence CHT (E2E, 1x64, CHT) outperforms multi-sequence batching (E2E, 4x16, CHE) with BN-free backbones but fails with BatchNorm. Partial freezing did not have a clear benefit and even underperformed with GN - likely due to fewer long-term dependencies. Despite not reaching SOTA performance (as the fully Transformer-based model by Chen et al. (2022) performs exceptionally well), our simple ConvNeXt and GN models consistently achieve the second highest scores, outperforming Transformer- and TCN-based multi-stage approaches (Trans-SVNet, TeCNO) as well as 3D-CNNs. BN-based models do not reach this level and again fail when using CHT, which was most effective for BN-free backbones.
@@ 4.3 Anticipation of Instrument Usage
Hypothesis 8: BatchNorm enables “cheating” in certain tasks. BatchNorm can leak information within batches to “cheat” certain objectives. While previous work finds empirical evidence for this effect through degrading performance (Wu and Johnson, 2021), we conclusively demonstrate it by carefully designing a toy task: We propose an impossible task which cannot be solved without “cheating”: Given training batches of size 2, predict for each input sample whether it is equal to the other sample in the batch. We use a small neural network with BatchNorm (Fig. 6, left) and with single binary inputs $x_{i}\in\{0,1\}$. Although this task is clearly unsolvable when each sample in the batch is processed individually, the BN model “cheats” by making predictions based on batch statistics (Fig. 6, right). Models expectedly fail (always predict $False$) when using global statistics during inference or networks without BN.
@@
Hypothesis 10: End-to-end models outperform state of the art in anticipation. Finally, we compare our models and strategies to previous work in instrument anticipation (Rivoir et al., 2020; Yuan et al., 2021). In Table 6, we observe that ConvNeXt and FrozenBN outperform previous work. The GN model only performs well when trained completely end-to-end but gives poor results with partial CNN freezing. As expected, BN models completely fail with single-sequence batches ($1\times 64,1\times 256$) but also provide subpar performance with multiple long sequences ($2\times 256,4\times 256,4\times 64$). Likely, this is in part caused by BN’s “cheating” (see Fig. 6(a)). However, other factors may have contributed since the partially frozen GN model underperforms as well. The end-to-end BN model trained on shorter 16-frame sequences performs surprisingly well. Likely, 16-second batches are too short for effective “cheating”.
@@
Anticipation error normal-$\downarrow$ (wMAE in sec.)
SOTA End-to-end AlexNet-LSTM (Rivoir et al., 2020) $66.02(\pm 0.85)$
2-stage ResNet-TCN (Yuan et al., 2021) (variant with only visual features) $65.34(\pm 0.59)$
@@ 5 Limitations and Scope
Our proposed strategies aim at increasing temporal context, and thus seem mostly applicable to tasks where this is relevant, e.g. long videos and workflow-based tasks such as cooking or other complex activities. Its impact on other video tasks remains open. Further, our experiments focus on 2D backbones, due to limitations in the surgical domain, and LSTMs, to show the effectiveness of simple, state-based end-to-end models for online tasks. While the incompatibility of BN and single-sequence batches is general to the video setting, investigating the feasibility and effectiveness of end-to-end approaches for other domains, backbones and temporal models requires further work and falls outside the scope of our study. Liu et al. (2022a) recently provided an extensive, insightful study on end-to-end temporal action detection with a wider range of models. However, BN issues are not discussed, although the common FrozenBN solution is utilized. Our work, on the other hand, aims at providing a deeper understanding of BN’s limitations in long-video settings but restricts practical implications to a smaller range of models.
@@ Appendix A Results for Surgical Phase Recognition
Table 7: Accuracy scores (mean + sample std. over runs) for surgical phase recognition on TeCNO’s (Czempiel et al., 2020) split (40 train. videos, 8 val., 32 test) used for our main analysis. Per run, accuracy is computed per video and then averaged. We highlight scores which outperform all 2-stage approaches by at least its sample standard deviation. Note that these results were presented in the form of figures in the main paper
Accuracy (video-wise) normal-$\uparrow$ Fig.
@@
TeCNO (Czempiel et al., 2020) (original) $88.56(\pm 0.27)$
@@
$8\times 8$, sliding eval. $85.9(\pm 0.1)$ $85.2(\pm 0.6)$ $83.0(\pm 2.2)$ $85.9(\pm 0.2)$ 2(a)
@@
$8\times 8$, carry-hidden eval. $89.4(\pm 0.4)$ $89.0(\pm 0.4)$ $86.8(\pm 1.3)$ $88.1(\pm 0.2)$ 2(b)
@@
$1\times 64$, carry-hidden train. $\color[rgb]{1,0,0}55.6(\pm 3.1)$ $89.4(\pm 0.4)$ $\bf 90.5(\pm 0.9)$ $\bf 92.3(\pm 1.5)$ 3 & 2(b)
@@
$4\times 64$, carry-hidden eval. $90.1(\pm 2.2)$ 2(c)
@@
$1\times 64$, carry-hidden train. - $\bf 92.4(\pm 0.4)$ $89.7(\pm 1.8)$ $\bf 93.3(\pm 0.3)$ 2(c)
@@
Table 8: F1 scores (mean + sample std. over runs) for surgical phase recognition on TeCNO’s (Czempiel et al., 2020) split (40 train. videos, 8 val., 32 test) used for our main analysis. Per training run, we compute the frame-based confusion matrix over the entire test set, compute a single F1 score per phase and average over phases. This strategy deviates from video-wise approaches in previous work but is less ambiguous and gives smoother scores in edge cases. We highlight scores which outperform all 2-stage approaches by at least its sample standard deviation
F1 (frame-wise) normal-$\uparrow$ Fig.
@@
$8\times 8$, sliding eval. $79.3(\pm 0.9)$ $78.8(\pm 0.2)$ $74.6(\pm 2.1)$ $78.4(\pm 1.2)$ 2(a)
@@
$8\times 8$, carry-hidden eval. $83.2(\pm 1.0)$ $82.8(\pm 0.8)$ $79.4(\pm 0.8)$ $81.9(\pm 0.6)$ 2(b)
@@
$1\times 64$, carry-hidden train. $\color[rgb]{1,0,0}40.0(\pm 4.1)$ $82.2(\pm 0.2)$ $83.9(\pm 0.7)$ $\bf 86.2(\pm 1.7)$ 3 & 2(b)
@@
$4\times 64$, carry-hidden eval. $84.5(\pm 1.4)$ 2(c)
@@
$1\times 64$, carry-hidden train. - $\bf 85.3(\pm 1.3)$ $82.4(\pm 1.0)$ $\bf 87.4(\pm 0.7)$ 2(c)
@@
Learning rate Accuracy (video-wise) normal-$\uparrow$
@@ Appendix B Results for Online Action Segmentation
Tables 10 and 11 show our results on the GTEA (Fathi et al., 2011) and 50Salads (Stein and McKenna, 2013) datasets. These are the same results which were presented as bar plots in Fig. 5 of the paper. Differently from most previous work, we formulate action segmentation as an online task on these datasets. We use the suggested splits for K-fold cross validation but for each training run, we use one set for validation (model selection), one set for testing and the remaining for training. We use the same hyperparameters as for surgical phase recognition (see Sec. E).
Table 10: Accuracy scores for online action segmentation on the GTEA dataset (Fathi et al., 2011). We use the suggested splits for 4-fold cross validation. Note that most works propose offline methods. For each backbone, we highlight the training strategy (CHE vs. CHT) which performs better. Note that these results were presented in the form of a bar plot in the main paper.
@@
Accuracy (video-wise) normal-$\uparrow$
@@
Table 11: Accuracy scores for online action segmentation on the 50Salads dataset (Stein and McKenna, 2013). We use the suggested splits for 5-fold cross validation. Note that most works propose offline methods. For each backbone, we highlight the training strategy (CHE vs. CHT) which performs better. Note that these results were presented in the form of a bar plot in the main paper.
@@
Accuracy (video-wise) normal-$\uparrow$
@@ Appendix C Model Architectures
Fig. 8 shows the network architectures used in our experiments. For the GroupNorm-based Resnet-50, we use GroupNorm with 32 groups since this performed best in the original paper (Wu and He, 2018) and has a pretrained model available [^3] https://github.com/ppwwyyxx/GroupNorm-reproduce/releases/tag/v0.1. For 2-stage TCN models, we follow TeCNO (Czempiel et al., 2020) as close as possible. Although not reported in the paper, we follow their exact model configuration with 2 stages, 9 layers per stage and 64 feature channels per layer [^4] https://github.com/tobiascz/TeCNO/issues/6#issuecomment-802851853.
@@
Fig. 8: Overview of model architectures used in our experiments. For instrument anticipation, we follow the original task formulation (Rivoir et al., 2020). For each of 5 instruments, 4 prediction values are required (1 for duration prediction and 3 for the auxiliary classification task)
@@ D.1 Surgical Phase Recognition
Surgical phase recognition on the Cholec80 dataset (Twinanda et al., 2016) is a dense temporal segmentation task with 7 phases and is therefore a simple multi-class classification when viewed at the frame level. Given a frame $x$ with corresponding class label $c\in\{1,\dots,7\}$ and a softmax-activated neural network $f$ which outputs a $7$D vector per input image, we use the standard cross-entropy loss for optimization:
@@ D.2 Anticipation of Surgical Instrument Usage
For instrument ancitipation we follow the task formulation from previous work (Rivoir et al., 2020). We anticipate 5 different instruments on the Cholec80 dataset, where the main task is to predict the remaining duration until occurrence of each instrument at each time point up to a horizon of 5 minutes. As an auxiliary classification task, we also predict whether each instrument is present and whether or not it will appear within the next 5 minutes (i.e. inside or outside of horizon).
@@
We then jointly optimize both tasks with the SmoothL1 (Yengera et al., 2018) loss for regression and cross entropy for classification:
@@ Appendix E Training Details & Hyperparameters
    Despite our use of different learning rates, our decision rule was very simple: We use a default $10^{-5}$ for models where all parameters are trained (End-to-end & Backbone only) and increased the learning by a factor of $10$ for shallower models (Partially frozen). For training only the temporal models in 2-stage approaches (Temporal only), we follow TeCNO (Czempiel et al., 2020) and use $5\cdot 10^{-4}$, which seemed to perform best in our experiments, too.
@@ E.1 Varying Hyperparameters on Other Datasets
Cholec80 (Twinanda et al., 2016) $1$ $216\times 384$ $7$ $80$
AutoLaparo (Wang et al., 2022) $1$ $216\times 384$ $7$ $21$
CATARACTS (Zisimopoulos et al., 2018) $5$ $216\times 384$ $19$ $50$
GTEA (Fathi et al., 2011) $15$ $216\times 288$ $11$ $28$
50Salads (Stein and McKenna, 2013) $5$ $216\times 288$ $19$ $50$
@@ Appendix F Metrics
Relaxed metrics for SOTA comparison: Some methods (including the current state-of-the-art methods Trans-SVNet (Gao et al., 2021) and TMRNet (Jin et al., 2021)) report relaxed versions of metrics on the CHolec80 dataset, which were proposed in the M2CAI 2016 Challenge (Stauder et al., 2016). Here, errors around phase transitions are ignored due to their ambiguity. Implementation details can be found in TMRNet’s code [^5] https://github.com/YuemingJin/TMRNet/blob/main/code/eval/result/matlab-eval/.
Weighted Mean Absolute Error (wMAE): We follow the original instrument anticipation approach (Rivoir et al., 2020) and report wMAE as our main metric. This is the mean of the frame-wise mean absolute errors inside (inMAE) and outside (outMAE) of the horizon. Specifically for each instrument, predicitons on all test videos are concatenated into a single sequence, to compute single inMAE and outMAE (and thus wMAE) scores per instrument, which are then averaged.

### v5

@@ 4.2 Surgical Phase Recognition
CATARACTS: Since CATARACTS contains shorter videos with more fine-grained annotations, the benefits of our context-maximizing strategies are expectedly smaller here. Nevertheless, we can still reproduce several findings (Table 5): Single-sequence CHT (E2E, 1x64, CHT) outperforms multi-sequence batching (E2E, 4x16, CHE) with BN-free backbones but fails with BatchNorm. Partial freezing did not have a clear benefit and even underperformed with GN - likely due to fewer long-term dependencies. Despite not reaching SOTA performance (as the fully Transformer-based model by Chen et al. [12] performs exceptionally well), our simple ConvNeXt and GN models consistently achieve the second highest scores, outperforming Transformer- and TCN-based multi-stage approaches (Trans-SVNet, TeCNO) as well as 3D-CNNs. BN-based models do not reach this level and again fail when using CHT, which was most effective for BN-free backbones.
@@ 4.3 Anticipation of Instrument Usage
Hypothesis 8: BatchNorm enables “cheating” in certain tasks. BatchNorm can leak information within batches to “cheat” certain objectives. While previous work finds empirical evidence for this effect through degrading performance [77], we conclusively demonstrate it by carefully designing a toy task: We propose an impossible task which cannot be solved without “cheating”: Given training batches of size 2, predict for each input sample whether it is equal to the other sample in the batch. We use a small neural network with BatchNorm (Fig. 6, left) and with single binary inputs $x_{i}\in\{0,1\}$. Although this task is clearly unsolvable when each sample in the batch is processed individually, the BN model “cheats” by making predictions based on batch statistics (Fig. 6, right). Models expectedly fail (always predict $False$) when using global statistics during inference or networks without BN.
@@
Hypothesis 10: End-to-end models outperform state of the art in anticipation. Finally, we compare our models and strategies to previous work in instrument anticipation [58, 86]. In Table 6, we observe that ConvNeXt and FrozenBN outperform previous work. The GN model only performs well when trained completely end-to-end but gives poor results with partial CNN freezing. As expected, BN models completely fail with single-sequence batches ($1\times 64,1\times 256$) but also provide subpar performance with multiple long sequences ($2\times 256,4\times 256,4\times 64$). Likely, this is in part caused by BN’s “cheating” (see Fig. 7(a)). However, other factors may have contributed since the partially frozen GN model underperforms as well. The end-to-end BN model trained on shorter 16-frame sequences performs surprisingly well. Likely, 16-second batches are too short for effective “cheating”.
@@
Anticipation error $\downarrow$ (wMAE in sec.)
SOTA End-to-end AlexNet-LSTM [58] $66.02(\pm 0.85)$
2-stage ResNet-TCN [86] (variant with only visual features) $65.34(\pm 0.59)$
@@ 5 Limitations and Scope
Our proposed strategies aim at increasing temporal context, and thus seem mostly applicable to tasks where this is relevant, e.g. long videos and workflow-based tasks such as cooking or other complex activities. Its impact on other video tasks remains open. Further, our experiments focus on 2D backbones, due to limitations in the surgical domain, and LSTMs, to show the effectiveness of simple, state-based end-to-end models for online tasks. While the incompatibility of BN and single-sequence batches is general to the video setting, investigating the feasibility and effectiveness of end-to-end approaches for other domains, backbones and temporal models requires further work and falls outside the scope of our study. Liu et al. [47] recently provided an extensive, insightful study on end-to-end temporal action detection with a wider range of models. However, BN issues are not discussed, although the common FrozenBN solution is utilized. Our work, on the other hand, aims at providing a deeper understanding of BN’s limitations in long-video settings but restricts practical implications to a smaller range of models.
@@ Appendix A Results for Surgical Phase Recognition
Table 7: Accuracy scores (mean + sample std. over runs) for surgical phase recognition on TeCNO’s [Czempiel et al., 2020] split (40 train. videos, 8 val., 32 test) used for our main analysis. Per run, accuracy is computed per video and then averaged. We highlight scores which outperform all 2-stage approaches by at least its sample standard deviation. Note that these results were presented in the form of figures in the main paper
Accuracy (video-wise) $\uparrow$ Fig.
@@
TeCNO [Czempiel et al., 2020] (original) $88.56(\pm 0.27)$
@@
$8\times 8$, sliding eval. $85.9(\pm 0.1)$ $85.2(\pm 0.6)$ $83.0(\pm 2.2)$ $85.9(\pm 0.2)$ 3(a)
@@
$8\times 8$, carry-hidden eval. $89.4(\pm 0.4)$ $89.0(\pm 0.4)$ $86.8(\pm 1.3)$ $88.1(\pm 0.2)$ 3(b)
@@
$1\times 64$, carry-hidden train. $\color[rgb]{1,0,0}55.6(\pm 3.1)$ $89.4(\pm 0.4)$ $\bf 90.5(\pm 0.9)$ $\bf 92.3(\pm 1.5)$ 3 & 3(b)
@@
$4\times 64$, carry-hidden eval. $90.1(\pm 2.2)$ 3(c)
@@
$1\times 64$, carry-hidden train. - $\bf 92.4(\pm 0.4)$ $89.7(\pm 1.8)$ $\bf 93.3(\pm 0.3)$ 3(c)
@@
Table 8: F1 scores (mean + sample std. over runs) for surgical phase recognition on TeCNO’s [Czempiel et al., 2020] split (40 train. videos, 8 val., 32 test) used for our main analysis. Per training run, we compute the frame-based confusion matrix over the entire test set, compute a single F1 score per phase and average over phases. This strategy deviates from video-wise approaches in previous work but is less ambiguous and gives smoother scores in edge cases. We highlight scores which outperform all 2-stage approaches by at least its sample standard deviation
F1 (frame-wise) $\uparrow$ Fig.
@@
$8\times 8$, sliding eval. $79.3(\pm 0.9)$ $78.8(\pm 0.2)$ $74.6(\pm 2.1)$ $78.4(\pm 1.2)$ 3(a)
@@
$8\times 8$, carry-hidden eval. $83.2(\pm 1.0)$ $82.8(\pm 0.8)$ $79.4(\pm 0.8)$ $81.9(\pm 0.6)$ 3(b)
@@
$1\times 64$, carry-hidden train. $\color[rgb]{1,0,0}40.0(\pm 4.1)$ $82.2(\pm 0.2)$ $83.9(\pm 0.7)$ $\bf 86.2(\pm 1.7)$ 3 & 3(b)
@@
$4\times 64$, carry-hidden eval. $84.5(\pm 1.4)$ 3(c)
@@
$1\times 64$, carry-hidden train. - $\bf 85.3(\pm 1.3)$ $82.4(\pm 1.0)$ $\bf 87.4(\pm 0.7)$ 3(c)
@@
Learning rate Accuracy (video-wise) $\uparrow$
@@ Appendix B Results for Online Action Segmentation
Tables 10 and 11 show our results on the GTEA [Fathi et al., 2011] and 50Salads [Stein and McKenna, 2013] datasets. These are the same results which were presented as bar plots in Fig. 5 of the paper. Differently from most previous work, we formulate action segmentation as an online task on these datasets. We use the suggested splits for K-fold cross validation but for each training run, we use one set for validation (model selection), one set for testing and the remaining for training. We use the same hyperparameters as for surgical phase recognition (see Sec. E).
Table 10: Accuracy scores for online action segmentation on the GTEA dataset [Fathi et al., 2011]. We use the suggested splits for 4-fold cross validation. Note that most works propose offline methods. For each backbone, we highlight the training strategy (CHE vs. CHT) which performs better. Note that these results were presented in the form of a bar plot in the main paper.
@@
Accuracy (video-wise) $\uparrow$
@@
Table 11: Accuracy scores for online action segmentation on the 50Salads dataset [Stein and McKenna, 2013]. We use the suggested splits for 5-fold cross validation. Note that most works propose offline methods. For each backbone, we highlight the training strategy (CHE vs. CHT) which performs better. Note that these results were presented in the form of a bar plot in the main paper.
@@
Accuracy (video-wise) $\uparrow$
@@ Appendix C Model Architectures
Fig. 8 shows the network architectures used in our experiments. For the GroupNorm-based Resnet-50, we use GroupNorm with 32 groups since this performed best in the original paper [Wu and He, 2018] and has a pretrained model available [^3] https://github.com/ppwwyyxx/GroupNorm-reproduce/releases/tag/v0.1. For 2-stage TCN models, we follow TeCNO [Czempiel et al., 2020] as close as possible. Although not reported in the paper, we follow their exact model configuration with 2 stages, 9 layers per stage and 64 feature channels per layer [^4] https://github.com/tobiascz/TeCNO/issues/6#issuecomment-802851853.
@@
Fig. 8: Overview of model architectures used in our experiments. For instrument anticipation, we follow the original task formulation [Rivoir et al., 2020]. For each of 5 instruments, 4 prediction values are required (1 for duration prediction and 3 for the auxiliary classification task)
@@ D.1 Surgical Phase Recognition
Surgical phase recognition on the Cholec80 dataset [Twinanda et al., 2016] is a dense temporal segmentation task with 7 phases and is therefore a simple multi-class classification when viewed at the frame level. Given a frame $x$ with corresponding class label $c\in\{1,\dots,7\}$ and a softmax-activated neural network $f$ which outputs a $7$D vector per input image, we use the standard cross-entropy loss for optimization:
@@ D.2 Anticipation of Surgical Instrument Usage
For instrument ancitipation we follow the task formulation from previous work [Rivoir et al., 2020]. We anticipate 5 different instruments on the Cholec80 dataset, where the main task is to predict the remaining duration until occurrence of each instrument at each time point up to a horizon of 5 minutes. As an auxiliary classification task, we also predict whether each instrument is present and whether or not it will appear within the next 5 minutes (i.e. inside or outside of horizon).
@@
We then jointly optimize both tasks with the SmoothL1 [Yengera et al., 2018] loss for regression and cross entropy for classification:
@@ Appendix E Training Details & Hyperparameters
    Despite our use of different learning rates, our decision rule was very simple: We use a default $10^{-5}$ for models where all parameters are trained (End-to-end & Backbone only) and increased the learning by a factor of $10$ for shallower models (Partially frozen). For training only the temporal models in 2-stage approaches (Temporal only), we follow TeCNO [Czempiel et al., 2020] and use $5\cdot 10^{-4}$, which seemed to perform best in our experiments, too.
@@ E.1 Varying Hyperparameters on Other Datasets
Cholec80 [Twinanda et al., 2016] $1$ $216\times 384$ $7$ $80$
AutoLaparo [Wang et al., 2022] $1$ $216\times 384$ $7$ $21$
CATARACTS [Zisimopoulos et al., 2018] $5$ $216\times 384$ $19$ $50$
GTEA [Fathi et al., 2011] $15$ $216\times 288$ $11$ $28$
50Salads [Stein and McKenna, 2013] $5$ $216\times 288$ $19$ $50$
@@ Appendix F Metrics
Relaxed metrics for SOTA comparison: Some methods (including the current state-of-the-art methods Trans-SVNet [Gao et al., 2021] and TMRNet [Jin et al., 2021]) report relaxed versions of metrics on the CHolec80 dataset, which were proposed in the M2CAI 2016 Challenge [Stauder et al., 2016]. Here, errors around phase transitions are ignored due to their ambiguity. Implementation details can be found in TMRNet’s code [^5] https://github.com/YuemingJin/TMRNet/blob/main/code/eval/result/matlab-eval/.
Weighted Mean Absolute Error (wMAE): We follow the original instrument anticipation approach [Rivoir et al., 2020] and report wMAE as our main metric. This is the mean of the frame-wise mean absolute errors inside (inMAE) and outside (outMAE) of the horizon. Specifically for each instrument, predicitons on all test videos are concatenated into a single sequence, to compute single inMAE and outMAE (and thus wMAE) scores per instrument, which are then averaged.